Silver Enterprises is planning to acquire Jurion Inc. Silver Enterprises has 8,300 shares outstanding with a share price of $46 and Jurion Inc. has 3,400 shares with a share price of $21. The synergy gains from the acquisition are expected to be $12,600

A. If Jurion Inc. is willing to be acquired for $24 per share in cash, what is the NPV of the merger? B. What is the price per share of the merged firm? C. What is the takeover premium?
Answer #1

A.NPV of mereger = Present value of benefits - cash paid

= 3,400*21 + 12,600 - 3,400*24

= $2,400

B. Price per share of merged firm = (8,300*46 + 3,400*21 + 12,600- 3,400*24)/8,300

= $46.29

Takeover premium = 24-21 = $3

or 3/21 = 14.29%